Listed below are pulse rates​ (beats per​ minute) from samples of adult males and females. Find the mean and median for each of the two samples and then compare the two sets of results. Does there appear to be a​ difference?


1
Expert's answer
2021-09-28T01:09:54-0400

Solution:

Let the missing data be



Mean of male"=\bar X=\dfrac{67+54+...+69}{15}=66.133"

Arranging in descending order: 88, 86, 77, 76, 69, 67, 64, 63, 62, 61, 59, 57, 55, 54, 54

Median = (n+1)/2 th term = 8 th term = 63

Mean of female"=\bar Y=\dfrac{90+79+...+74}{15}=81.066"

Arranging in descending order: 93, 90, 89, 89, 88, 87,87, 87, 84, 79, 74, 72, 68, 66,63.

Median = (n+1)/2 th term = 8 th term = 87


The mean and median for males are lower than that of females.

Thus, the pulse rate for females appear to be higher than males.